# The Best Instruction-Tuning Data are Those That Fit

### v2

@@ 5.5 Further Experiments
We extend our evaluation using OpenHermes-2.5 (Teknium, 2023), a high-quality, open instruction-tuning dataset containing approximately 1 million distinct instructions. Following a similar strategy to the previous combo experiment, we source responses from additional datasets from  Huang et al. (2024) and HuggingFace-H4 (2024). For preference-based datasets, we include only the winning responses, consistent with the methodology in the Tulu-Olmo experiment. This process results in 575K unique instructions and 1.34 million total instruction-response pairs. As shown in Table 3, GRAPE-selected data yields better performances too. The consistent improvement across models reaffirm our earlier observations: GRAPE serves as an adaptive response selection mechanism that significantly enhances SFT performance that has wide applicability.
5.6 Discussion:GRAPE works if all responses are from the same LM
GRAPE
Qwen2.5
-72B
Llama3.1
-405B
Gemma
-it-9B
LC 28.1 25.8 16.3 26.9
WR 33.1 24.3 17.0 20.2
Table 4: Alpaca-Eval2 Results On Magpie-Zoo (Xu et al., 2024d).
In Section 4, we demonstrated how GRAPE enables practitioners to refine model-generated responses for improved training outcomes. Beyond that, GRAPE can optimize responses from a single generator. Using the Magpie-Zoo instruction set (Xu et al., 2024d), we sample 10 responses per instruction from Qwen2.5-72B-Instruct, select in-distribution responses with GRAPE, to train a Mistral-v0.3-7B model. We also compare with training on top-3 reported datasets of Magpie-Zoo. As shown in Table 4, GRAPE-selected responses further boost the performance. Practically, batch sampling of responses cause little latency (Zhong et al., 2024; Zhou et al., 2024d). One therefore can sample batches of responses from a single strong model and apply GRAPE on top of it to find the response closest to the base model’s distribution efficiently.
5.7 Discussion: Is In-Distribution A Silver Bullet?
The success of GRAPE verifies our key hypothesis that matching SFT distribution to the base model benefits the performance. Yet, we argue that this cannot be pushed to the limit of SFT using data from the same distribution. By experimenting with training on solutions sampled from a trained version of the same base model, we notice a drastic performance drop from the original model as shown in Table 5.
Model Data Avg.
Self-Distilled 28.4(-)
Mistral-7B Original-UI 32.7
Self-Distilled 29.4(-)
Llama3.1-8B Original-UI 36.8
Self-Distilled 15.1(-)
Llama3.2-3B Original-UI 20.3
Table 5: Performance Degradation From Self-Distillation On UI.
We explored this using the MATH (Hendrycks et al., 2021b) dataset. A base model was fine-tuned using solutions from stronger models like Llama3.1-70B-Instruct or GPT-3.5-Turbo-augmented MetaMathQA (Yu et al., 2024). This fine-tuned model then generated new solutions, which were used to further fine-tune another base model instance, simulating iterative on-policy tuning.
Dataset Model Response Generator N Acc.
MATH Mistral Llama3.1-70B 10 18.2
MATH Mistral FT-Mistral 10 15.9 (-)
MATH Llemma Llama3.1-70B 10 26.2
MATH Llemma FT-Llemma 10 23.6 (-)
MATH Mistral MM-AnsAug - 22.3
MATH Mistral FT-Mistral 10 20.6 (-)
MATH Llemma MM-AnsAug - 28.1
MATH Llemma FT-Llemma 10 21.4 (-)
Table 6: Effect of self-generation on MATH dataset. FT-Mistral refers to the model right above that row finetuned from either MM-AnsAug or Llama3.1-70B-Instruct produced solutions. N stands for the number of responses sampled.
As shown in Table 6, performance declined when solutions were sampled from the fine-tuned model itself. This decline stems from reduced solution diversity, leading to distributional drift and poorer generalization.
These results validate our claim in §2 that robust fine-tuning requires more than alignment with the base model—it needs complementary strategies to maintain response diversity and quality.
6 Related Works
Data Engineering For Instruction Tuning
Data is central to the success of effective instruction tuning,  (Xu et al., 2023; Xia et al., 2024; Chan et al., 2024), featuring both automated data synthesis (Xu et al., 2024a; Zeng et al., 2024; Yu et al., 2024; Wei et al., 2023) and selection (Xia et al., 2024; Chen et al., 2023a; Parkar et al., 2024; Li et al., 2024e). Some selection approaches focus on high-quality data by leveraging LLMs (Chen et al., 2023a; Parkar et al., 2024; Li et al., 2024b) or employing principled metrics (Kang et al., 2024; Mekala et al., 2024; Xia et al., 2024), while others, such as  Yang et al. (2024c); Das & Khetan (2023), aim to identify diversity-optimized subsets for greater efficiency.
Another emerging trend is the customization of training data based on the characteristics of the base models. For instance,  Li et al. (2024c); Du et al. (2023) leverage the base model itself to select a subset of instructions, while  Li et al. (2024a) introduce a teacher model to guide the selection process. However, these approaches primarily focus on identifying a re-weighted subset of questions for training. In contrast, GRAPE has a different focus, where it aims to find a set of responses from the solution space that not only provide good coverage of the golden distribution but also align closely with the base model’s policy.
On-Policy Methods For Language Model Alignment
Recent advances in RLHF (Ouyang et al., 2022), RLAIF (Lee et al., 2024b) and preference learning (Rafailov et al., 2023) research have identified the benefits of on-policy data for model alignment (Xu et al., 2024b; Tajwar et al., 2024; Zhang et al., 2024c; Guo et al., 2024b; Liu et al., 2024d; Zhang et al., 2024a; Miao et al., 2024; Gulcehre et al., 2023). However, these works primarily emphasize the preference-tuning stage, where models are refined to generate samples from an already reasonable policy. In contrast, less attention is given to the supervised fine-tuning phase, which starts with base models that are not yet well-prepared to generate high-quality answers that align sufficiently with the desired policy for optimization.
7 Conclusion
We present GRAPE, a simple yet highly effective approach to improve supervised fine-tuning data building on our empirically-verified hypothesis that instruction tuning data shall better match the base model’s distribution to optimize the training outcome. GRAPE is surprisingly simple and efficient, where it customizes response data to supervise each model by taking highest-probability instances. We demonstrated the effectiveness of GRAPE on multiple settings. Remarkably, GRAPE can outperform not only SOTA data selection baselines, but also Tulu3-SFT by using a 1/3 subset of its SFT instructions, and performs better than training on 4.5 times data it selects from. We show GRAPE as a highly scalable and simple approach can be adopted in multiple practical scenarios with minimal overhead.
Appendix A Parameter Distance
We measure the L2 norm of model parameter difference between fine-tuned and pre-trained checkpoints, as a signal of how much the distribution has drifted during SFT (ichi Amari, 2016; Cover & Thomas, 2006). We notice that training over well-matched distribution shifts the parameter less than training over those ill-matched.
Mistral-7B-v0.3 Llama3.1-8B Qwen2.5
GRAPE 8.006 8.196 8.426
Worst 8.029 8.202 8.467
Table 7: Performance comparison across different models
Appendix B Further Details On S2L
@@
This section details the experimental setup for two unsupervised data selection baseline, S2L applied to the joined pool of Tulu-3 and Olmo-v2 data.
@@
Appendix C Further Training Details
@@
Appendix D Further Ablations on UltraInteract.
@@
Appendix E Additional Related Works On Model Dependent Data Selection Approaches
E.1 0. Notations
@@ Appendix E Additional Related Works On Model Dependent Data Selection Approaches
1. TLDR: The Final Table
@@
GRAPE (ours) 00 $NF_{\theta}$
@@
In-run gradient-based influence $C(\theta,D,1)$ 00
@@
Embedding-based 00 $NF_{\theta}$
Simple uncertainty indicators 00 $NF_{\theta}$
@@
2. Gradient-based Methods
@@
2.1 Gradient-based Influence
Gradient-based influence computes the pairwise influence scores between each pair of training and validation examples. Training data with the highest influence are selected, as training on them leads to the theoretically largest decrease in model loss on validation data. LESS (Xia et al., 2024) formulates the pairwise influence scores as the cosine similarity between the gradients of training and validation data, and computes these gradient features using the following two steps:
@@ 2.1 Gradient-based Influence
2.2 Gradient Matching
Gradient matching also requires per-sample gradients, but utilizes their information in a different way. It performs clustering based on these gradient features to group similar data, and then applies an iterative greedy selection algorithm. In order to scale to LLM-level gradient computation and clustering, TAGCOS (TAGCOS) completely follows the warmup training and gradient computation pipeline of LESS (LESS). As the computational bottleneck here is still the gradient computation instead of clustering or iterative selection, (TAGCOS) also shares the same computational cost as (Xia et al., 2024):
@@ 2.2 Gradient Matching
2.3 Gradient Norm
The $L_{2}$-norms of gradient vectors can also serve as effective indicators for data selection. (Paul et al., 2023) proposes GraNd, which obtains a utility score for each training point based on its gradient norm early in the training. More specifically, it starts from $m$ different model weight initializations, trains each model on the whole dataset to obtain per-sample gradient norms, and finally averages the $m$ gradient norms for each training point to obtain the final GraNd score. Therefore, the computational cost of GraNd is shown below:
@@ 2.3 Gradient Norm
3. Embedding-based Methods
@@
Under a supervised data selection setup where validation data representing target task distributions are available, Representation-based Data Selection (RDS; (Rubin et al., 2022; Hanawa et al., 2021)) computes the embedding similarity between training and validation data, and selects training points that are most similar to the target distribution in the embedding space.
For an unsupervised setup where only the embeddings of training data are accessible, geometry-based coreset sampling methods are widely used (Qin et al., 2024). Grounded on the intuition that close samples in the embedding space often share similar properties, a diverse subset can be obtained by controlling the minimum distance between any two selected data points. Among them, using K-center greedy sampling to select embedding-based facility locations has been proven especially effective for instruction fine-tuning of LLMs (Bhatt et al., 2024a).
@@
4. LogProb-based Methods
@@ 4.1 Simple Uncertainty-based Indicators
Some simple model-based indicators inspired by the notion of uncertainty have been shown effective for a long time and recently extended to data selection for LLM instruction tuning (Marion et al., 2023b; Bhatt et al., 2024b). (Bhatt et al., 2024b) demonstrates the effectiveness of various indicators including mean entropy, least confidence, mean margin, etc. These simple indicators do not require additional training and can also be directly obtained with the pretrained final model $\theta$. Their computational cost is shown below:
@@ 4.2 Perplexity (PPL)
A common approach is to use the final model $\theta$ as the reference model $\theta_{\text{ref}}$ to ensure the alignment in PPL patterns (Marion et al., 2023a), but prior work (Ankner et al., 2024) also shows that a reference model much smaller than the final model can also be an effective PPL-based data selector. The computational cost for PPL-based selection is shown below:
@@ 4.3 Learnability
4.4 Loss Trajectory
Moreover, logprob-based methods can also obtain finer-grained information from the training dynamics of LLMs. S2L (Yang et al., 2024b) obtains a feature vector for each training point by collecting their training loss trajectories over $T$-epoch training on a small reference model $\theta_{\text{ref}}$, and then applies K-means clustering to equally sample data points from each trajectory cluster. Prior work shows its superiority over other logprob-based indicators, but it also comes with significant computational cost:
  • Additional training: $C(\theta_{\text{ref}},D,T)$. Here the choice of $\theta_{\text{ref}}$ is especially important, as prior work (Xia et al., 2023) shows that reference models that come from the same model family as the final model tend to have similar loss trajectories of training data, so they can preserve more fidelity in their loss trajectory patterns.

### v3

@@ B.2 Experiment on OpenHermes
To test the generality of our findings beyond the UltraInteract and Tulu-Olmo settings, we conduct additional experiments on the OpenHermes-2.5 (Teknium, 2023) dataset—a large-scale, high-quality instruction-tuning corpus with approximately 1 million distinct instructions.
Following the setup from §5, we apply GRAPE to select from responses aggregated across sources, including Huang et al. (2024) and HuggingFace-H4 (2024). For preference-based datasets, we retain only the winning responses to ensure quality, mirroring our earlier selection protocol. This results in 575K unique instructions and 1.34M instruction-response pairs.
As shown in Table 8, GRAPE continues to outperform naive combination strategies. The consistent gains across diverse data sources and model families strengthen our central claim: GRAPE is a general-purpose, model-aligned response selection strategy that reliably improves SFT performance in real-world, large-scale instruction tuning.
B.3 Data Selection For Long Chain-of-Thoughts
O1-/R1-style long chain-of-thoughts have drawn increasing attention. This paradigm, exemplified by models like OpenAI’s O1 and DeepSeek-R1, has shown remarkable success in challenging domains such as mathematics and coding. We further experiment with the use of GRAPE in long chain-of-thought distillation. We generate multiple candidate trajectories using R1-Distill-Qwen-1.5B DeepSeek-AI et al. (2025) for a subset of OpenR1-Math (Face, 2025) dataset — LUFFY (Yan et al., 2025), and verify the correctness of each, retaining the correct ones.
We compare the results on lowest versus highest perplexity instances below in table 9.
Model Perplexity Acc.
Qwen2.5-1.5B Highest 0.330
Qwen2.5-1.5B Lowest (GRAPE) 0.396
Qwen2.5-3B Highest 0.524
Qwen2.5-3B Lowest (GRAPE) 0.544
Table 9: Performance metrics on MATH dataset
B.4 Token-level GRAPE
To further investigate the alternative uses of our insight, we conduct experiments beyond data selection by incorporating token-level likelihoods directly into the training objective. Specifically, we modified the loss function to weigh each token proportionally to its likelihood. We trained this variant on the OpenThoughts‑114k Team (2025) dataset, a curated collection of 114k high-quality reasoning samples spanning domains such as math, science, code etc. Evaluation followed the same benchmark suite as LUFFY Yan et al. (2025), including competition-level math datasets (AIME24/25, AMC, MATH‑500, Minerva, OlympiadBench) and general reasoning tests (ARC‑c, GPQA‑Diamond, MMLU‑Pro). Our result in Table 10 show that our token-level likelihood-weighted training yields consistent improvements across several benchmarks.
Benchmark Baseline (%) Token-level GRAPE (%)
MATH 55.60 60.20
OlympiadBench 21.04 28.30
Minerva 15.07 19.85
AIME-24 3.12 4.90
AMC 23.76 29.74
AIME-25 4.38 3.23
ARC-C 60.84 76.62
GPQA-Diamond 7.07 19.70
MMLU-Pro 27.02 34.35
Table 10: Qwen2.5-3B performance (in %) on various benchmarks under Baseline vs. Token-level GRAPE.
Appendix C Further Details On Baselines
@@
This section details the experimental setup for our data selection baselines: S2L,LESS and NV-Embed.
C.1 S2L
@@
C.2 LESS
LESS is a state-of-the-art model-based and supervised data selection method that leverages gradient-based influence estimation. Given a small set of validation examples per task, LESS computes the influence of each training example by measuring the weighted cosine similarity between their LoRA gradients across multiple warmup checkpoints. It then aggregates these influence scores by averaging over validation examples within each task, followed by taking the maximum across tasks to obtain a scalar utility score per training example. Training examples are selected greedily based on these scores. In our experiments, we use the same base model for both selection and training, and follow the original LESS setup: 5% warmup training for 4 epochs and a gradient projection dimension of 8192.
C.3 NV-Embed
Embedding-based data selection as detailed in (Ivison et al., 2025) is a supervised data selection method that ranks training examples by computing cosine similarity between their embeddings and those of validation examples. Unlike model-aware methods like LESS, embedding-based data selection is model-agnostic: it relies on fixed, pretrained embedding models (in our case, we used NV-embed-v2, the state-of-the-art embedding model) rather than the target model. Instead of aggregating similarity scores into a single utility value per training example, embedding-based data selection uses a round-robin strategy that iteratively selects the highest-scoring example for each validation instance, ensuring diverse coverage across tasks. We follow the original setup from Ivison et al. (2025) in our experiments.
Appendix D Further Training Details
@@
Appendix E Further Ablations on UltraInteract.
See Tables 11 and 4.
@@
Appendix F Additional Related Works On Model Dependent Data Selection Approaches
Model-dependent data selection methods leverage internal signals from a target model—such as gradients, embeddings, or log-probabilities—to identify training examples that are most useful for fine-tuning. These approaches have led to strong empirical results across various settings. However, many of them involve substantial computational costs, such as repeated gradient computations or auxiliary model training, which can limit their scalability. We discuss these approaches and the costs they incur in this section.
F.1 Notations
@@
F.2 TLDR: The Final Table
@@
GRAPE (ours) 0 $NF_{\theta}$
@@
In-run gradient-based influence $C(\theta,D,1)$ 0
@@
Embedding-based 0 $NF_{\theta}$
Simple uncertainty indicators 0 $NF_{\theta}$
@@
F.3 Gradient-based Methods
@@
F.3.1 Gradient-based Influence
Gradient-based influence computes the pairwise influence scores between each pair of training and validation examples. Training data with the highest influence are selected, as training on them leads to the theoretically largest decrease in model loss on validation data. LESS Xia et al. (2024) formulates the pairwise influence scores as the cosine similarity between the gradients of training and validation data, and computes these gradient features using the following two steps:
@@ F.3.1 Gradient-based Influence
F.3.2 Gradient Matching
Gradient matching also requires per-sample gradients, but utilizes their information in a different way. It performs clustering based on these gradient features to group similar data, and then applies an iterative greedy selection algorithm. In order to scale to LLM-level gradient computation and clustering, TAGCOS Zhang et al. (2024b) completely follows the warmup training and gradient computation pipeline of LESS Xia et al. (2024). As the computational bottleneck here is still the gradient computation instead of clustering or iterative selection, Zhang et al. (2024b) also shares the same computational cost as Xia et al. (2024):
@@ F.3.2 Gradient Matching
F.3.3 Gradient Norm
The $L_{2}$-norms of gradient vectors can also serve as effective indicators for data selection. Paul et al. (2023) proposes GraNd, which obtains a utility score for each training point based on its gradient norm early in the training. More specifically, it starts from $m$ different model weight initializations, trains each model on the whole dataset to obtain per-sample gradient norms, and finally averages the $m$ gradient norms for each training point to obtain the final GraNd score. Therefore, the computational cost of GraNd is shown below:
@@ F.3.3 Gradient Norm
F.4 Embedding-based Methods
@@
Under a supervised data selection setup where validation data representing target task distributions are available, Representation-based Data Selection (RDS; Rubin et al. (2022); Hanawa et al. (2021)) computes the embedding similarity between training and validation data, and selects training points that are most similar to the target distribution in the embedding space.
For an unsupervised setup where only the embeddings of training data are accessible, geometry-based coreset sampling methods are widely used Qin et al. (2024). Grounded on the intuition that close samples in the embedding space often share similar properties, a diverse subset can be obtained by controlling the minimum distance between any two selected data points. Among them, using K-center greedy sampling to select embedding-based facility locations has been proven especially effective for instruction fine-tuning of LLMs Bhatt et al. (2024).
@@
F.5 LogProb-based Methods
@@ 4.1 Simple Uncertainty-based Indicators
Some simple model-based indicators inspired by the notion of uncertainty have been shown effective for a long time and recently extended to data selection for LLM instruction tuning (Marion et al., 2023b; Bhatt et al., 2024). Bhatt et al. (2024) demonstrates the effectiveness of various indicators including mean entropy, least confidence, mean margin, etc. These simple indicators do not require additional training and can also be directly obtained with the pretrained final model $\theta$. Their computational cost is shown below:
@@ F.5.1 Perplexity (PPL)
A common approach is to use the final model $\theta$ as the reference model $\theta_{\text{ref}}$ to ensure the alignment in PPL patterns Marion et al. (2023a), but prior work Ankner et al. (2024) also shows that a reference model much smaller than the final model can also be an effective PPL-based data selector. The computational cost for PPL-based selection is shown below:
@@ F.5.2 Learnability
F.5.3 Loss Trajectory
Moreover, logprob-based methods can also obtain finer-grained information from the training dynamics of LLMs. S2L Yang et al. (2024b) obtains a feature vector for each training point by collecting their training loss trajectories over $T$-epoch training on a small reference model $\theta_{\text{ref}}$, and then applies K-means clustering to equally sample data points from each trajectory cluster. Prior work shows its superiority over other logprob-based indicators, but it also comes with significant computational cost:
  • Additional training: $C(\theta_{\text{ref}},D,T)$. Here the choice of $\theta_{\text{ref}}$ is especially important, as prior work Xia et al. (2023) shows that reference models that come from the same model family as the final model tend to have similar loss trajectories of training data, so they can preserve more fidelity in their loss trajectory patterns.
@@ F.5.3 Loss Trajectory
Appendix G Multi-Round GRAPE
To evaluate whether GRAPE continues to provide benefits after an initial round of fine-tuning, we conducted a second-round experiment using LLaMA3.1-8B. In the first round, we selected the best responses per instruction using GRAPE and fine-tuned the model accordingly. Then, we re-applied GRAPE on the newly fine-tuned model to select a fresh set of responses and conducted another round of fine-tuning. This second iteration led to further performance improvements across multiple benchmarks, including AlpacaEval, WizardEval (both original and LeetCode), and real-world tasks such as MATH, MMLU, and BBH. Notably, GRAPE Round 2 improved average benchmark scores from 35.9% to 37.3%, demonstrating that the method remains effective and even compounding when iteratively applied.
Method AE WR WR (LC) MATH MMLU BBH
GRAPE Round 1 15.2 14.8 19.4 32.1 64.5 69.6
GRAPE Round 2 17.6 19.6 18.9 33.2 64.5 70.0
Table 14: Performance of LLaMA3.1-8B after two rounds of GRAPE fine-tuning.
Appendix H Details Of Correlation Analysis
To analyze how well different models align with the training distribution, we conducted a perplexity-based study on the Tulu-v3 training set. Specifically, we randomly sampled 1,000 instances from the training data and used a wide array of generator models to produce responses for each instruction. For each response, we computed its perplexity under each base model and ranked the responses by perplexity, with lower perplexity indicating better matching of distribution. We then identified the top-1 response per instance for each base model and visualized the overall rankings using a heatmap (Figure 4).